# Modality Translation for Object Detection Adaptation Without Forgetting Prior Knowledge

### v1

@@ 4.1 Experimental Methodology
LLVIP: The LLVIP dataset is a surveillance dataset composed of $30,976$ images, in which $24,050$ ($12,025$ IR and $12,025$ RGB paired images) are used for training and $6,926$ for testing ($3,463$ IR and $3,463$ RGB paired images) with only pedestrians annotated. FLIR ALIGNED: For the FLIR, we used the sanitized and aligned paired sets provided by Zhang et al. [42], which has $10,284$ images, that is $8,258$ for training ($4,129$ IRs and $4,129$ RGBs) and $2,026$ ($1,013$ IRs and $1,013$ RGBs) for test. The FLIR images are taken from the perspective of a camera in the front of a car, and the resolution is $640$ by $512$. It contains classes of bicycles, dogs, cars, and people. It has been found that for the case of FLIR, the "dog" objects are inadequate for training [3], thus we decided to remove the dogs.
@@
In our experiments, we used $80\%$ of the training set for training and the rest for validation. All results reported are on the test set. For the FLIR dataset, we removed the dogs as there are not many annotations. As starting pre-trained weights for the detectors, we used Torchvision models with COCO [24] weights and for the U-Net translation network, we used PyTorch Segmentation Models [16] and we changed the last layer for 3-channel (RGB-like) with sigmoid, to be closer to an image with values between 00 and $1$, to perform translation instead of traditional segmentation. For the translation network backbones, we explored our default ResNet34, and for subsequent studies on reducing parameters, we dive into ResNet and MobileNet-family. All the code is available on GitHub for reproducibility in the experiments. To ensure fairness, we trained the detectors under the library version and the same experimental design, i.e., data order, augmentations, etc. Furthermore, we trained with PyTorch Lightning [8] training framework, evaluated the APs with TorchMetrics [7], and logged all experiments with WandB [1] logging tool. The different measured APs can be found in the supplementary material as additional metrics provided in this work.
@@ 4.2 Comparison with Translation Approaches
In this section, we compare the ModTr method with different image-to-image methods employing different learning strategies. These include basic image processing strategies[12], reconstruction strategies such as CycleGAN [45], CUT [32], and FastCUT [32], which employs a constitutive learning approach, as well as HalluciDet [25], which utilizes a detection-based loss. As outlined in Table 5, we evaluated the methods based on their final detection performance across three commonly used detectors: FCOS, RetinaNet, and FasterRCNN. The reported results are derived from the infrared test set and are averaged over three different seeds, which helps mitigate the impact of randomness across runs and splits of the training and validation datasets.
For each method, we also consider its dependency on the source data (RGB) and ground truth bounding boxes on the IR images (Box). Methods that rely on reconstruction techniques do not require box annotations on IR images but cannot provide accurate translations for detection purposes. However, HalluciDet and ModTr require box annotations to adjust the input image in a discriminative manner. The main difference between HalluciDet and ModTr is the use of source images. HalluciDet requires RGB images for an initial fine-tuning of the model, while our approach can work without that fine-tuning by reusing the detector’s zero-shot knowledge.
The proposed ModTr method demonstrates robustness across the three detectors and consistently exhibits improvement on two different datasets: LLVIP [18] and FLIR aligned [11]. It’s worth noting that each algorithm described in Table 5 employs different training supervisions. For instance, CycleGAN employs an adversarial mechanism with both RGB and infrared modalities in an unpaired setting. Similarly, CUT and FastCUT operate with positive and negative patches in an unpaired setting. In contrast, HalluciDet doesn’t require the presence of both modalities during training but employs a detection mechanism during training similar to ours. Note that in our approach, we solely require examples from the target modality. In this section, we present the performance of our best approach $\text{ModTr}_{\odot}$. For additional results, refer to the supplementary materials.
Table 1: Comparison of the detection performance (AP) of three different detectors (FCOS, RetinaNet, and FasterRCNN) over different image-to-image methods to translate the infrared to RGB-like images and our proposed method (ModTr). The methods were evaluated on the infrared test set of LLVIP and FLIR datasets. The column RGB tells whether the method needs access to RGB images during training, and Box refers to the use of ground truth bounding boxes during training.
@@
FCOS RetinaNet FasterRCNN
@@
Image translation RGB Box Test Set IR (Dataset: LLVIP)
FCOS RetinaNet FasterRCNN
@@
As depicted in Table 5, the detection performance of ModTr over the LLVIP dataset exhibited significant improvements. Specifically, it surpassed HalluciDet, the second best, by more than $29.0$ AP with both FCOS and RetinaNet architectures, while obtaining comparable results with FasterRCNN. Such disparity with the previous technique can be attributed to the loss of previous knowledge inherent in HalluciDet, which necessitates a pre-fine-tuning strategy on the source modality. Although the performance of the FLIR dataset also improved, the dataset’s inherent challenges, such as changing the background from a moving car setup, make detection more difficult. Nonetheless, our proposal consistently enhances results, with improvements of more than $11$ AP for FCOS and RetinaNet, and over $7$ AP for FasterRCNN. We also observed improvements on the AP50 and AP75. Because of the space constraint, we include these in Supplementary Materials. These promising results indicate that our proposal can effectively translate images from the original IR modality to an RGB-like representation, sufficiently close to the source data to be usable by the detector.
4.3 Fine-Tuning of the Translator and Detector
In the following ablation study, we further demonstrate that the proposed approach can be trained jointly with both translation and detector networks in a fully fine-tuning setting, which does not preserve the detector’s knowledge but significantly improves performance. Here, we compare the performance of the three proposed fusion functions with that of the standard detector’s fine-tuning.
As depicted in Tab. 2, for each fusion function we include variations with and without fine-tuning. The table shows AP for the LLVIP dataset, with a consistent trend across all detectors (FCOS, RetinaNet, and FasterRCNN). Furthermore, in the case of the FLIR dataset, we observed enhancements over the standard detector fine-tuning with the various fusion techniques, both with and without fine-tuning versions. As demonstrated, our approach surpasses standard fine-tuning while maintaining the detector’s performance in the original modality. It is worth noting that our method also improves performance in terms of localization metrics such as AP50 and AP75 compared to fine-tuning alone, and we provide detailed results in the Supplementary Materials.
Table 2: Comparison of the fine-tuning of the detector, the different non-parametric fusion strategies, and the fine-tuning version of the fusion. For each ModTr version, we have a correspondent fine-tuning train to update both the detector and ModTr network during training. The detection performance of three different detectors (FCOS, RetinaNet, and FasterRCNN) was evaluated on the infrared test set of LLVIP and FLIR datasets.
@@ 4.3 Fine-Tuning of the Translator and Detector
FCOS RetinaNet FasterRCNN
@@
ModTr+ 56.44 ± 0.75 53.18 ± 1.03 57.14 ± 0.50
ModTr 57.01 ± 0.71 54.43 ± 0.35 56.95 ± 0.37
ModTr 57.63 ± 0.66 54.83 ± 0.61 57.97 ± 0.85
@@
FCOS RetinaNet FasterRCNN
@@
ModTr+ 34.63 ± 0.24 33.70 ± 0.59 37.09 ± 0.74
ModTr 34.94 ± 0.52 33.72 ± 0.22 37.16 ± 0.47
ModTr 34.60 ± 0.38 33.85 ± 0.34 37.01 ± 0.15
@@ 4.4 Different Backbones for ModTr
In this context, we focused on evaluating the ModTr architecture and examining the trade-off between performance and parameter cost. It is widely recognized that increasing the number of parameters can enhance performance, but this relationship is not strictly linear. We demonstrated that models with fewer parameters can still achieve good performance; for example, MobileNetv2, with fewer parameters than ResNet18, sometimes outperformed it. This trade-off highlights the versatility of the model, which can be deployed with MobileNet-based architectures and utilized in low-cost devices. In Table 3, we successfully reduced the default number of parameters from $24.4$M (ResNet34) to $6.6$M using MobileNetv2 while maintaining similar performance. For instance, on LLVIP, MobileNetv2 achieved a mean AP of $56.15$, comparable to $56.35$ AP50 from ResNet34 (others APs and detectors are reported in the supplementary material).
This approach opens up new possibilities, particularly in scenarios where using one translation network and one detector (e.g., one ModTr and one detector for thermal/RGB) proves advantageous. This setup requires a total of $44.9$M parameters, compared to $83.6$M parameters, when employing two detectors—one for each modality (for example, for FasterRCNN). Similar reductions in parameter costs were observed for FCOS (from $66.4$M to $36.3$M) and RetinaNet (from $68$M to $37.1$M) when using one detector for both modalities while preserving the knowledge of the previous modality and incorporating a new one. These numbers are based on MobileNetv3s, which strikes a balance between performance and the number of parameters, making it suitable for memory-restricted systems. The complete evaluations for FCOS and RetinaNet are included in the supplementary material.
Table 3: Study on the translation backbone describing the tradeoff between the number of parameters for the translation network and the detection performance for each detector. The results comprise the three default detectors used (FCOS, RetinaNet, and FasterRCNN) in AP under different MobileNet and ResNet backbones for the translation network.
@@
$\text{ModTr}_{\oplus}$ Refer to caption
Figure 3: Illustration of a sequence of $8$ images of LLVIP and FLIR dataset for Faster R-CNN. For each dataset, the first row is the RGB modality, followed by the IR modality, followed by FastCUT, and different representations created by ModTr and their variations. For visualizations for other detectors, see Suppl. Material.
@@ 4.5 Knowledge Preservation through Input Modality Translation
In this section, we show how various adaptation paradigms, depicted in Figure 2, are capable of solving the final task while preserving intrinsic knowledge. We compare our proposed method, ModTr, to two other fine-tuning baseline methods. The first baseline method consists of N detectors, each of which fine-tunes the target modality individually. The second baseline method is a single detector trained on the joint modality with balanced sampling.
Table 4 illustrates the final performance. While both N detectors and single detectors exhibit relatively similar final performance, ModTr stands out as the only method capable of preserving knowledge while excelling at the final task. Specifically, both fine-tuning baselines compromised the detector’s zero-shot capability. This consistent trend is observable across various detectors and datasets.
Table 4: Average Precision (AP) comparison of knowledge preserving techniques N-Detectors, 1-Detector, and N-ModTr-1-Detector. The results comprise three detectors (FCOS, RetinaNet, and FasterRCNN) across diverse datasets such as LLVIP, FLIR, and COCO.
@@
COCO 00.22 ± 0.02 00.29 ± 0.01 35.48 ± 0.00
FasterRCNN LLVIP 59.62 ± 1.23 62.50 ± 1.29 57.97 ± 0.85
@@ 4.6 Visualization of ModTr Translated Images
In Figure 3, we present qualitative results for LLVIP and IR, alongside comparison with fine-tuning. Panel a) displays the ground-truth RGB, while b) showcases the results of fine-tuning using IR. Subsequently, in c), we present ModTr with addition-based fusion, in d) ModTr with attention-based fusion, and in e) ModTr with a Hadamard product-based fusion for each dataset over the Faster R-CNN detector. Due to space constraints, we provide additional visualizations for the other detectors and competitive methods in the supplementary material. Notably, the IR exhibits some false positives, especially when there is an overlap of the detected objects, which is mitigated by our method. It’s worth noting that, as described, the product-based fusion yields a darker image, contrasting with the additive characteristics of other fusion strategies. Further insights into the visualization are also provided in supplementary material, revealing how our method effectively blurs or removes objects that do not have elements of the target classes, thereby facilitating detection for the OD. Although the obtained intermediate representations are not visually pleasant, they prove more efficient for incorporating the knowledge necessary for the OD. Additionally, we conducted experiments with loss function terms aimed at enhancing the visual effects of the image, but they were not conclusive in terms of helping the detection performance.
4.7 Fine-tuning ModTr and the detector
The main reason to use ModTr is to avoid fine-tuning the detector for a specific task so that it can preserve its knowledge and be used for multiple modalities. However, in this section, we consider what would happen if we learn jointly ModTr and the detector weights. Results are reported in Figure 4. We see that fine-tuning the detector can further boost performance. Thus, another application of ModTr could be to use it to improve the fine-tuning of a detector with a reduced additional computational cost.
@@ 5 Conclusion
In this work, we present a novel ModTr method for adapting ODs without changing their parameters. The proposed approach performs well in different settings and results better than powerful image-to-image models and previous competitors. The proposed ModTr method was evaluated for different tasks, such as detection based on image translation, comparison with traditional fine-tuning, and incremental modality application. Experimental results showed the high level of performance and versatility of our method. Our approach also benefits from preserving the full knowledge of the detector, which opens the possibility of using the translation network as a node to change the modality for an unaltered detector. This is much more convenient in terms of flexibility and computation than having a specialized OD for each modality. Future works will expand the knowledge of the OD by incorporating additional text embedding to help the ModTr perform few-shot and zero-shot learning in the image translation space targeting detection.
Acknowledgments: This work was supported by Distech Controls Inc., the Natural Sciences and Engineering Research Council of Canada, the Digital Research Alliance of Canada, and MITACS.

### v3

@@ 4.1 Experimental Methodology
LLVIP: LLVIP is a surveillance dataset composed of $30,976$ images, in which $24,050$ ($12,025$ IR and $12,025$ RGB paired images) are used for training and $6,926$ for testing ($3,463$ IR and $3,463$ RGB paired images) with only pedestrians annotated. FLIR ALIGNED: We used the sanitized and aligned paired sets provided by Zhang et al. [50]. It has $10,284$ images, that is $8,258$ for training ($4,129$ IRs and $4,129$ RGBs) and $2,026$ ($1,013$ IRs and $1,013$ RGBs) for test. FLIR images are captured from the perspective of a camera in the front of a car, with a resolution of $640$ by $512$. It contains the bicycles, dogs, cars, and people classes. It has been found that with FLIR, the "dog" objects are inadequate for training [4], thus we decided to remove them.
@@
In our experiments, we randomly selected $80\%$ of the training set for training and the rest for validation. All results reported are on the test set. As starting pre-trained weights for the detectors, we used Torchvision models with COCO [30] weights and for the U-Net translation network, we used PyTorch Segmentation Models [21] and we changed the last layer for 3-channel (RGB-like) with a Sigmoid function, to be closer to an image with values between 00 and $1$, to perform translation instead of traditional segmentation. For the translation network backbones, we explored our default ResNet34, and for subsequent studies on reducing parameters, we dive into ResNet and MobileNet-family. All the code is available on GitHub for reproducibility in the experiments. To ensure fairness, we trained the detectors under the library version and the same experimental design, i.e., data order, augmentations, etc. Furthermore, we trained with PyTorch Lightning [10] training framework, evaluated the APs with TorchMetrics [8], and logged all experiments with WandB [2] logging tool. The different performance measures (e.g., APs) can be found in suppl. materials.
@@ 4.2 Comparison with Translation Approaches
In this section, ModTr is compared with different image-to-image translation methods employing different learning strategies. These include basic image processing strategy[15], reconstruction strategies such as CycleGAN [53], CUT [39], and FastCUT [39], which employs a contrastive learning approach, as well as HalluciDet [31], which utilizes a detection-based loss. As outlined in Table 1, we evaluated the methods based on their final detection performance across three commonly used detectors: FCOS, RetinaNet, and Faster R-CNN. The reported results are derived from the IR test set and are averaged over three different seeds, which helps mitigate the impact of randomness across runs and splits of the training and validation datasets.
For each method, we also consider its dependency on the prior knowledge data (RGB) and ground truth bounding boxes (bboxes) on the IR images. Methods that rely on reconstruction techniques do not require bbox annotations on IR images but cannot provide accurate translations for detection purposes. However, HalluciDet and ModTr require bbox annotations to adjust the input image in a discriminative manner. The main difference between HalluciDet and ModTr is the use of source images. HalluciDet requires RGB images for an initial fine-tuning of the model, while our approach can work without that fine-tuning by reusing the detector’s zero-shot knowledge.
The proposed ModTr displays robustness across the three detectors and consistently exhibits improvement on two different datasets: LLVIP [23] and FLIR aligned [13]. Note that each algorithm described in Table 1 employs different training supervisions. For instance, CycleGAN employs an adversarial mechanism with both RGB and infrared modalities in an unpaired setting. Similarly, CUT and FastCUT operate with positive and negative patches in an unpaired setting. In contrast, HalluciDet doesn’t require the presence of both modalities during training but employs a detection mechanism during training similar to ours. In our approach, we solely require examples from the target modality. In this section, we present the performance of our best approach $\text{ModTr}_{\odot}$. For additional results, refer to suppl. materials.
Table 1: Detection performance (AP) of ModTr versus baseline image-to-image methods to translate the IR to RGB-like images, using three different detectors (FCOS, RetinaNet, and Faster R-CNN). The methods were evaluated on IR test set of LLVIP and FLIR datasets. The RGB column indicates if the method required access to RGB images during training, and Box refers to the use of ground truth boxes during training.
@@
FCOS RetinaNet Faster R-CNN
@@
Image translation RGB Box Test Set IR (Dataset: FLIR)
FCOS RetinaNet Faster R-CNN
@@
As reported in Table 1, the detection performance of ModTr over the LLVIP dataset exhibited significant improvements. Specifically, it surpassed HalluciDet, the second best, by more than $29.0$ AP with both FCOS and RetinaNet architectures, while obtaining comparable results with Faster R-CNN. Such disparity with the previous technique can be attributed to the loss of previous knowledge inherent in HalluciDet, which necessitates a pre-fine-tuning strategy on the source modality. Although the performance of the FLIR dataset also improved, the dataset’s inherent challenges, such as changing the background from a moving car setup, make detection more difficult. Nonetheless, our proposal consistently enhances results, with improvements of more than $11$ AP for FCOS and RetinaNet, and over $7$ AP for Faster R-CNN. We also observed improvements on the AP50 and AP75. Because of the space constraint, we include these in supplementary materials. These promising results indicate that our proposal can effectively translate images from the original IR modality to an RGB-like representation, sufficiently close to the source data to be usable by the detector.
4.3 Translation vs. Fine-tuning
In this section, we further show that the proposed approach can be trained jointly with both translation and detector, which preserves the detector’s knowledge. Here, ModTr is compared to three baselines fully fine-tuning (FT), FT of the head and LoRA [19], and our best ModTr fusion strategy, as shown in  Tab. 2.
We conduct LoRA fine-tuning using two settings. In the first, we apply LoRA across all layers; in the second, only to the last layer of detectors. The latter results in superior performance, so we have adopted it as our default LoRA setting. The  Tab. 2 shows AP for the LLVIP and FLIR datasets, with a consistent trend across all detectors (FCOS, RetinaNet, and Faster R-CNN). Furthermore, in the case of the FLIR dataset, we observed enhancements of ModTr over the standard detector FT. As demonstrated, our approach surpasses standard fine-tuning while maintaining the detector’s performance in the original modality. It is worth noting that our method also improves performance in terms of localization metrics such as AP50 and AP75 compared to fine-tuning alone, and we provide detailed results in the supplementary materials.
Table 2: Detection performance (AP) of ModTr versus baseline fine-tuning (FT) of the detector, FT of the head and LoRA [19], using three different detectors (FCOS, RetinaNet, and Faster R-CNN. The methods were evaluated on IR test set of LLVIP and FLIR datasets. Results with "-" diverged from the optimization.
@@
FCOS RetinaNet Faster R-CNN
@@
FT Head 49.11 ± 0.70 44.00 ± 0.28 59.33 ± 2.17
LoRA [19] 47.72 ± 0.58 - 54.83 ± 1.30
ModTr (ours) 57.63 ± 0.66 54.83 ± 0.61 57.97 ± 0.85
@@
FCOS RetinaNet Faster R-CNN
@@
FT Head 27.40 ± 0.12 26.78 ± 0.70 33.53 ± 0.36
LoRA [19] - - 29.44 ± 0.61
ModTr (ours) 35.49 ± 0.94 34.27 ± 0.27 37.21 ± 0.46
@@ 4.4 Different Backbones for ModTr
In this context, we evaluate ModTr and examine the trade-off between performance and parameter cost. It is widely recognized that increasing the number of parameters can enhance performance, but this relationship is not strictly linear. We demonstrated that models with fewer parameters can still achieve good performance; for example, MobileNetv2, with fewer parameters than ResNet18, sometimes outperformed it. This trade-off highlights the versatility of the model, which can be deployed with MobileNet-based architectures and utilized in low-cost devices. In Table 3, the default number of parameters is successfully reduced from $24.4$M (ResNet34) to $6.6$M using MobileNetv2 while maintaining similar performance. For instance, on LLVIP, MobileNetv2 achieved a mean AP of $56.15$, comparable to $56.35$ AP50 from ResNet34 (others APs and detectors are reported in the supplementary material).
This approach opens up new possibilities, particularly in scenarios where using one translation network and one detector (e.g., one ModTr and one detector for RGB/IR) proves advantageous. This setup requires a total of $44.9$M parameters, compared to $83.6$M parameters, when employing two detectors—one for each modality (for example, for Faster R-CNN). Similar reductions in parameter costs were observed for FCOS (from $66.4$M to $36.3$M) and RetinaNet (from $68$M to $37.1$M) when using one detector for both modalities while preserving the knowledge of the previous modality and incorporating a new one. These numbers are based on MobileNetv3s, which strikes a balance between performance and the number of parameters, making it suitable for memory-restricted systems. The complete evaluations for FCOS and RetinaNet are included in the supplementary material.
Table 3: Detection performance (AP) of ModTr with different backbones for the translation networks with different numbers of parameters, using three different detectors (FCOS, RetinaNet, and Faster R-CNN). The methods were evaluated on IR test set of LLVIP and FLIR datasets.
@@
Figure 3: Illustration of a sequence of $8$ images of LLVIP and FLIR dataset for Faster R-CNN. For each dataset, the first row is the RGB modality, followed by the IR modality and different representations created by ModTr. For visualizations of other detectors and variants of ModTr, please refer to the supplementary materials.
@@ 4.5 Knowledge Preservation through Input Modality Translation
ModTr is designed to prevent catastrophic forgetting by keeping the weights of the pre-trained detector fixed. In this section, we demonstrate how various adaptation paradigms, shown in Figure 2, effectively solve the final task while preserving intrinsic knowledge. We compare our proposed method, ModTr, with two fine-tuning baseline methods. The first baseline method involves N-detectors, each fine-tuning the target modality individually. The second baseline method employs a single detector trained on the joint modality using balanced sampling. Note that while a copy of the original detector can be used in the N-detectors paradigm, it is unavailable in the $1$-detector paradigm because the original modality is assumed to be inaccessible during training.
In all scenarios, we use COCO as the pre-training dataset and LLVIP and FLIR as target domains. Specifically, in the N-detectors scenario (Fig.2a), we fine-tune one detector on each dataset and use a copy of the original detector for the RGB modality. In the $1$-detector scenario (Fig.2c), we fine-tune one detector on the combined FLIR and LLVIP datasets. In the N-ModTr-1-Detector scenario (Fig.2b), two translators are trained, one per dataset. To assess catastrophic forgetting, we re-evaluate each scenario on COCO-val.
Table 4 shows the final performance. While all adaptation paradigms achieve relatively similar performance, the $1$-detector method completely fails in the zero-shot scenario. The N-detectors method mitigates this by duplicating the detector three times. In contrast, ModTr preserves knowledge using a single detector and three efficient translators, demonstrating its practicality for embedded devices, as it requires less memory. Based on the average performance on all datasets, ModTr obtains the best results.
Table 4: Detection performance (AP) of knowledge preserving techniques N-Detectors, 1-Detector, and N-ModTr-1-Detector, using three different detectors (FCOS, RetinaNet, and Faster R-CNN). The methods were evaluated on COCO and IR test sets of LLVIP and FLIR datasets.
@@
COCO 35.48 ± 0.00 00.29 ± 0.01 35.48 ± 0.00
AVG. 39.24 ± 0.76 26.24 ± 1.28 41.52 ± 0.29
Faster R-CNN LLVIP 59.62 ± 1.23 62.50 ± 1.29 57.97 ± 0.85
@@ 4.6 Visualization of ModTr Translated Images
In Figure 3, we present qualitative results for LLVIP and FLIR, alongside a comparison with fine-tuning. Each dataset section includes three rows: the first row displays the ground-truth RGB images, the second row showcases the results of fine-tuning using IR, and the last row features ModTr with a Hadamard product-based fusion over the Faster R-CNN detector. Due to space constraints, additional visualizations for other detectors and fusion strategies are provided in the supplementary materials. Notably, the IR results exhibit some false positives, particularly when detected objects overlap. Our method mitigates this issue effectively. Further insights, provided in the supplementary materials, reveal how our method effectively blurs or removes objects that do not belong to the target classes, thereby enhancing detection accuracy. Although the obtained intermediate representations are not visually pleasant, they prove more efficient for incorporating the knowledge necessary for the OD. Additionally, we conducted experiments with loss function terms aimed at enhancing the visual effects of the image, but they were not conclusive in terms of helping the detection performance.
4.7 Fine-tuning of ModTr and the Detector
The main reason to use ModTr is to avoid fine-tuning the detector for a specific task so that it can preserve its knowledge and be used for multiple modalities. However, in this section, we consider what would happen if we learn jointly ModTr and the detector weights. Results are reported in Figure 4. We see that fine-tuning the detector can further boost performance. Thus, another application of ModTr could be used to improve the fine-tuning of a detector with a reduced additional computational cost.
@@ 5 Conclusion
In this paper, a novel method called ModTr is proposed for adapting RGB object detectors (ODs) for IR modality without changing their parameters. A key advantage of our approach is that it preserves the full knowledge of the detector, allowing the translation network to act as a node that changes the modality for an unaltered detector. This is much more flexible and computationally efficient than having a specialized OD for each modality. Our approach performs well in various settings, outperforming powerful image-to-image models and previous competitors. We evaluated ModTr for different tasks, including detection based on image translation, comparison with traditional fine-tuning, and incremental IR modality application. Experimental results show the high performance and versatility of our method in all these settings.
Additionally, to explore integrating modalities beyond IR, we applied ModTr to Canny edges extracted from IR images as detailed in the supplementary material. While ModTr significantly enhances the performance of zero-shot RGB OD on edges, it still does not match the effectiveness of full fine-tuning on this other modality. We believe this shortfall arises from the limited information provided by edges compared to the richer data in the IR modality, leading to lower initial zero-shot OD performance. A potential solution is to replace the deterministic translator module within ModTr with a generative model. This substitution could enrich modality information by generating the missing data, potentially improving the zero-shot detector’s performance. This promising direction will be explored in future research.
Acknowledgments
This work was supported in part by Distech Controls Inc., the Natural Sciences and Engineering Research Council of Canada, the Digital Research Alliance of Canada, and MITACS.